(4x^4 - 48)^(1/4) = x
A 2
B{-2,2}
C{-3,4}
D4
E {-4,-2,2}
F no solution
x/3SQRT(2X-3) - 1/SQRT(2x-3) = 1/3

Answer by lwsshak3(11628)

(4x^4 - 48)^(1/4) = x
raise both sides to the 4th power
4x^4-48=x^4
3x^4=48
x^4=16
x=2 (ans A)